# Phase space analysis of CCDM cosmologies

### v1

@@ III.5.2 Two fluid system
Creation rate Critical point Existence Stability $\Omega_{dm}$ Acceleration
$\Gamma_{0}+\Gamma_{1}H$ $F^{\prime}_{1}(1,0)$ Always Stable if $\beta\geq 3$ and Unstable if $\beta<3$ $1$ $q=(1-\beta)/2$, yes if $\beta>1$
$F^{\prime}_{2}\left(1,\frac{3-\beta}{3-\beta+\alpha}\right)$ $\beta\leq 3$ Stable $1$ $q=-1$, always yes
$\Gamma_{0}+\Gamma_{3}H^{-1}$ $G^{\prime}_{1}(1,0)$ Always Unstable $1$ $q=1/2$, no
$G^{\prime}_{2}\left(1,\frac{6}{6+\alpha+\sqrt{\alpha^{2}+12\mu}}\right)$ Always Stable $1$ $q=-1$, always yes
$\Gamma_{0}+\Gamma_{2}H^{2}$ $I^{\prime}_{\pm}\left(1,\frac{3+2\gamma\pm\sqrt{9-4\alpha\gamma}}{2(3+\alpha+\gamma)}\right)$ $4\alpha\gamma\leq 9$ $I^{\prime}_{+}:$ Stable and $I^{\prime}_{-}:$ Unstable $1$ $q=-1$, always yes
$\Gamma_{1}H+\Gamma_{2}H^{2}$ $J^{\prime}_{1}(1,1)$ Always Unstable if $\beta\geq 3$ and Stable if $\beta<3$ $1$ $q=(1-\beta)/2$, yes if $\beta>1$
$J^{\prime}_{2}\left(1,\frac{\gamma}{\gamma+3-\beta}\right)$ $\beta\leq 3$ Unstable $1$ $q=-1$, always yes
$\Gamma_{1}H+\Gamma_{3}H^{-1}$ $L^{\prime}_{1}(1,0)$ Always Unstable if $\beta<3$ and Stable if $\beta\geq 3$ $1$ $q=(1-\beta)/2$, yes if $\beta>1$
$L^{\prime}_{2}\left(1,\frac{3-\beta-\sqrt{(3-\beta)\mu}}{3-\beta-\mu}\right)$ $\beta\leq 3,~{}\beta+\mu\neq 3$ Stable $1$ $q=-1$, always yes
$\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}$ $M^{\prime}_{\pm}\left(1,\frac{3-\beta+2\gamma\pm\sqrt{(3-\beta)^{2}-4\alpha\gamma}}{2(3+\alpha-\beta+\gamma)}\right)$ $(3-\beta)^{2}-4\alpha\gamma\geq 0$, $M^{\prime}_{+}:$ Stable $1$ $q=-1$, always yes
$3+2\alpha>\beta$, and $M^{\prime}_{-}:$ Unstable
$3+2\gamma>\beta$
Table 6: Properties of the critical points of the dynamical system (43), (47), (50), (53), (57) and (60) for the matter creation rate $\Gamma=\Gamma_{0}+\Gamma_{1}H,~{}\Gamma_{0}+\Gamma_{3}H^{-1},~{}\Gamma_{0}+\Gamma_{2}H^{2},~{}\Gamma_{1}H+\Gamma_{2}H^{2},~{}\Gamma_{1}H+\Gamma_{3}H^{-1}$ and $\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}$ respectively.
Two-fluid system

### v2

@@ III.6 Model: $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$
Parameter ($\mathbf{\xi}$) Critical point Existence Eigenvalues Stability $\bf{\Omega_{\rm dm}^{\rm eff}}$ $\bf{\Omega_{f}}$ $\mathbf{q}$ Acceleration
$N_{1}\left(0,0\right)$ Always $\left(\frac{3(1+w-\xi)}{\xi},\frac{3(1+w)}{2}\right)$ Stable if $\xi>1+w$, w<1w<-1; 0 11 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Unstable if $\xi<1+w$;
Saddle if $\xi>1+w$, w>1w>-1
$N_{2}\left(1,0\right)$ Always $\left(-3(1+w-\xi),\frac{3\xi}{2}\right)$ Saddle if $\xi<1+w$; 11 0 $\frac{3\xi}{2}-1$ $\xi<\frac{2}{3}$
Unstable if $\xi>1+w$
$N_{3}\left(0,1\right)$ Always $\left(\frac{3(1+w-\xi)}{\xi},-\frac{3(1+w)}{2}\right)$ Stable if $\xi>1+w$, w>1w>-1; 0 11 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Otherwise Saddle
$\xi>0$ $N_{4}\left(1,1\right)$ Always $\left(-3(1+w-\xi),-\frac{3\xi}{2}\right)$ Stable if $\xi<1+w$; 11 0 $\frac{3\xi}{2}-1$ $\xi<\frac{2}{3}$
Saddle if $\xi>1+w$
$N_{5}\left(0,z_{c}\right)$ w=1w=-1 (3,0)(-3,0) Stable 0 11 1-1 Yes
$N_{6}\left(x_{c},0\right)$ $\xi=1+w$ $\left(\frac{3(1+w)}{2},0\right)$ Unstable $x_{c}$ $1-x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
$N_{7}\left(x_{c},1\right)$ $\xi=1+w$ $\left(-\frac{3(1+w)}{2},0\right)$ Stable $x_{c}$ $1-x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
$P_{1}\left(0,0\right)$ $0\leq x<\frac{\xi}{\xi-1}$ $\left(\frac{3(1+w-\xi)}{\xi-1},\frac{3(1+w)\xi}{2(\xi-1)}\right)$ Stable if $\xi<1+w$, w<1w<-1; 0 11 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Otherwise Saddle
$P_{2}\left(1,0\right)$ $\frac{\xi}{\xi-1}<x\leq 1$ $\left(\frac{3(1+w-\xi)}{\xi-1},\frac{3\xi}{2(1-\xi)}\right)$ Stable if $\xi<1+w$; 11 0 $\frac{3\xi}{2}-1$ $\xi<\frac{2}{3}$
Saddle if $\xi>1+w$
$P_{3}\left(0,1\right)$ $0\leq x<\frac{\xi}{\xi-1}$ $\left(\frac{3(1+w-\xi)}{\xi-1},-\frac{3(1+w)\xi}{2(\xi-1)}\right)$ Stable if $\xi<1+w$, w>1w>-1; 0 11 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Unstable if $\xi>1+w$, w<1w<-1;
Otherwise Saddle
$\xi<0$ $P_{4}\left(1,1\right)$ $\frac{\xi}{\xi-1}<x\leq 1$ $\left(\frac{3(1+w-\xi)}{\xi-1},\frac{3\xi}{2(\xi-1)}\right)$ Saddle if $\xi<1+w$; 11 0 $\frac{3\xi}{2}-1$ $\xi<\frac{2}{3}$
Unstable if $\xi>1+w$
$P_{5}\left(0,z_{c}\right)$ $0\leq x\leq\frac{\xi}{\xi-1}$, w=1w=-1 $\left(\frac{3\xi}{1-\xi},0\right)$ Stable 0 11 1-1 Yes
$P_{6}\left(x_{c},0\right)$ $\xi=1+w$, $x_{c}\neq\frac{\xi}{\xi-1}$ $(\phi_{\mp},0)$ Stable $x_{c}$ $1-x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
$P_{7}\left(x_{c},1\right)$ $\xi=1+w$, $x_{c}\neq\frac{\xi}{\xi-1}$ $(\psi_{\pm},0)$ Unstable $x_{c}$ $1-x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Table 6: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical systems (31), (32) and (33) with the matter creation rate $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$. Here, the eigenvalues $\phi_{\mp}$, $\psi_{\pm}$ are given by $\phi_{\mp}=\mp\frac{3(1+w)\left(1+w\left(1-x_{c}\right)\right)}{2w}$ and $\psi_{\pm}=\pm\frac{3(1+w)\left(1+w\left(1-x_{c}\right)\right)}{2w}$. Specifically, $\phi_{-}$ and $\psi_{+}$ are relevant in the region $\frac{\xi}{\xi-1}<x\leq 1$, while $\phi_{+}$ and $\psi_{-}$ correspond to $0\leq x<\frac{\xi}{\xi-1}$.
Figure 16: Description of the phase space controlled by the matter creation rate $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$. Upper Left Plot: The phase plot of the system (31) when we have assumed $w=0.1$ and $\xi=0.4$. Upper Middle Plot: The phase plot of the system (31) considering w=0.4w=-0.4 and $\xi=0.7$. Upper Right Plot: The phase space of the system (31) when the EoS $w$ takes the value 1-1. Here we use $\beta=1$ but any positive value of $\beta$ gives same type of phase portrait. Lower Left Plot: The phase plot of the system (31) when we assume w=1.4w=-1.4 and $\xi=1.4$. Lower Middle Plot: The phase plot of the systems (32) and (33) when we assume $w=0.1$ and $\xi=-0.7$. Lower Right Plot: The phase portrait of the dynamical systems (32) and (33) when we take w=1.4w=-1.4 and $\xi=-0.8$. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (i.e. q<1q<-1). The red line corresponding to the singular line $x=\frac{\xi}{\xi-1}$ of the system (31) separates the domain ${\bf R}=[0,1]^{2}$ into two parts.
Figure 17: The figure illustrates the evolution of key cosmological parameters namely, the dark matter density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the decelerating parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ for the dynamical system (31) with matter creation rate $\Gamma=2\left(1-\frac{1}{\xi}\right)\frac{\dot{H}}{H}$. Left Plot: We have drawn this plot using the model parameters $w=0.1$ and $\xi=0.6$. Here, the trajectories end in DM dominated accelerated quintessence era. As no decelerating DM dominated saddle point emerges (see the upper left plot of Fig. 16), a past DM-dominated phase is not achievable in this case. Middle Plot: This plot is produced adopting the values of the model parameters w=0.6w=-0.6 and $\xi=1.2$ where the late time evolution of the universe is attracted by DE dominated quintessence era connecting through a DM dominated decelerated era. Right Plot: For this plot, we take w=1.2w=-1.2 and $\xi=1.2$ where one can see that trajectories are attracted by DE dominated phantom era at late-times joining through a DM dominated decelerated phase.
[…]